Suppose Zoea Inc is contemplating investing in a new business. An existing company in the target industry has an equity beta of 0.80 but is very conservatively financed with a market value, equity-to-value ratio 95%. Zoea intends to use an equity-to-value ratio of 40% and wonders what beta to use in estimating a cost of capital in the new business.
Answer #1

First we need to find out the unlevered beta

Unlevered beta = beta / ( 1 + debt / equity )

= 0.80 / ( 1 + (1-0.95)/0.95 ) = 0.76

levered beta = unlevered beta * ( 1 + debt / equity) = 0.76 * ( 1 + (1-0.4)/0.4) = 1.90

beta to use in estimating a cost of capital in the new business = 1.90